Flexible Overhead Budget Leno Manufacturing Company prepared the following factory overhead cost budget for the Press Department for October of the current year, during which it expected to require 12,000 hours of productive capacity in the department: Variable overhead cost: Indirect factory labor $99,600 Power and light 5,640 Indirect materials 36,000 Total variable overhead cost $141,240 Fixed overhead cost: Supervisory salaries $49,430 Depreciation of plant and equipment 31,070 Insurance and property taxes 19,770 Total fixed overhead cost 100,270 Total factory overhead cost $241,510 Assuming that the estimated costs for November are the same as for October, prepare a flexible factory overhead cost budget for the Press Department for November for 10,000, 12,000, and 14,000 hours of production. Round your interim computations to the nearest cent, if required. Enter all amounts as positive numbers. Leno Manufacturing Company Factory Overhead Cost Budget-Press Department For the Month Ended November 30 Direct labor hours 10,000 12,000 14,000 Variable overhead cost: Indirect factory labor $ $ $ Power and light Indirect materials Total variable factory overhead $ $ $ Fixed factory overhead cost: Supervisory salaries $ $ $ Depreciation of plant and equipment Insurance and property taxes Total fixed factory overhead $ $ $ Total factory overhead cost $ $ $
Answer #1
Leno Manufacturing Company
Factory Overhead Cost Budget-Press Department
For the Month Ended November 30
Direct labor hours 10000 12000 14000
Variable overhead cost:
Indirect factory labor 83000 99600 116200
Power and light 4700 5640 6580
Indirect materials 30000 36000 42000
Total variable factory overhead 117700 141240 164780
Fixed factory overhead cost:
Supervisory salaries 49430 49430 49430
Depreciation of plant and equipment 31070 31070 31070
Insurance and property taxes 19770 19770 19770
Total fixed factory overhead 100270 100270 100270
Total factory overhead cost 217970 241510 265050
Workings:
Direct labor hours 10000 14000
Variable overhead cost:
Indirect factory labor =99600/12000*10000 =99600/12000*14000
Power and light =5640/12000*10000 =5640/12000*14000
Indirect materials =36000/12000*10000 =36000/12000*14000